# Multimodal Fusion with Relational Learning for Molecular Property Prediction

### v1

@@ Abstract
Graph-based molecular representation learning is essential for accurately predicting molecular properties in drug discovery and materials science; however, it faces significant challenges due to the intricate relationships among molecules and the limited chemical knowledge utilized during training. While contrastive learning is often employed to handle molecular relationships, its reliance on binary metrics is insufficient for capturing the complexity of these interactions. Multimodal fusion has gained attention for property reasoning, but previous work has explored only a limited range of modalities, and the optimal stages for fusing different modalities in molecular property tasks remain underexplored. In this paper, we introduce MMFRL (Multimodal Fusion with Relational Learning for Molecular Property Prediction), a novel framework designed to overcome these limitations. Our method enhances embedding initialization through multi-modal pre-training using relational learning. We also conduct a systematic investigation into the impact of modality fusion at different stages—early, intermediate, and late—highlighting their advantages and shortcomings. Extensive experiments on MoleculeNet benchmarks demonstrate that MMFRL significantly outperforms existing methods. Furthermore, MMFRL enables task-specific optimizations. Additionally, the explainability of MMFRL provides valuable chemical insights, emphasizing its potential to enhance real-world drug discovery applications.
@@ 1 Introduction
Graph representation learning for molecules has gained significant attention in drug discovery and materials science, as it effectively encapsulates molecular structures and enables the effective investigation of structure-activity relationships (Wieder et al., 2020; Zhang et al., 2022; Fang et al., 2022; Wang et al., 2023). In this paradigm, atoms are treated as nodes and chemical bonds as edges, effectively encapsulating the connectivities that define molecular behavior. However, it poses significant challenges due to intricate relationships among molecules and the limited chemical knowledge utilized during training.
@@
Figure 1: Multimodal Fusion with Relational Learning for Molecular Property Prediction (MMFRL). This figure shows our proposed idea about how to transfer the knowledge from other modalities and use fusion to improve the performance further. Unlike the general contrastive learning framework shown in Appendix Figure A.2, MMFRL doesn’t need to define positive or negative pairs and is capable of learning continuous ordering from target similarity.
@@
Often, contrastive learning (CL) is employed to study relationships among molecules, but it relies on binary metrics of positive and negative pairs, and tends to oversimplify complex molecular interactions. For example, consider Thalidomide: while the (R)- and (S)-enantiomers share the same topological graph and differ only at a single chiral center, their biological activities are drastically different—the (R)-enantiomer is effective in treating morning sickness, whereas the (S)-enantiomer causes severe birth defects. In other words, the (R)- and (S)-enantiomers are similar in terms of topological stucture but dissimilar in terms of biological activities. Thus, a more sophisticated approach is required to tackle these scenarios. A potential solution would be to use continuous metrics within a multi-view space, enabling a more comprehensive understanding of these complex molecular relationships.
@@
When it comes to multimodal learning for molecules, we often encounter data availability and incompleteness issues. This raises a critical question: how can multimodal information be effectively leveraged for molecular property reasoning when such data is absent in downstream tasks? Recent studies have demonstrated the effectiveness of pretraining molecular Graph Neural Networks (GNNs) by integrating additional knowledge sources (Wang et al., 2021; 2022b; Liu et al., 2022a; Xu et al., 2023a). Building on this foundation, a promising solution is to pretrain multiple replicas of molecular GNNs, with each replica dedicated to learning from a specific modality. This approach allows downstream tasks to benefit from multimodal data that is not accessible during fine-tuning, ultimately improving representation learning.
@@
2 Preliminaries

### v2

@@ Abstract
Graph-based molecular representation learning is essential for predicting molecular properties in drug discovery and materials science. Despite its importance, current approaches struggle with capturing the intricate molecular relationships and often rely on limited chemical knowledge during training. Multimodal fusion, which integrates information from molecular graph and other data modalities, has emerged as a promising avenue for enhancing molecular property prediction. However, existing studies have explored only a narrow range of modalities, and the optimal integration stages for multimodal fusion remain largely unexplored. Furthermore, the reliance on auxiliary modalities poses challenges, as such data is often unavailable in downstream tasks. Here, we present MMFRL (Multimodal Fusion with Relational Learning), a framework designed to address these limitations by leveraging relational learning to enrich embedding initialization during multimodal pre-training. MMFRL enables downstream models to benefit from auxiliary modalities, even when these are absent during inference. We also systematically investigate modality fusion at early, intermediate, and late stages, elucidating their unique advantages and trade-offs. Using the MoleculeNet benchmarks, we demonstrate that MMFRL significantly outperforms existing methods with superior accuracy and robustness. Beyond predictive performance, MMFRL enhances explainability, offering valuable insights into chemical properties and highlighting its potential to transform real-world applications in drug discovery and materials science.
@@ 1 Introduction
Graph representation learning for molecules has gained significant attention in drug discovery and materials science, as it effectively encapsulates molecular structures and enables the effective investigation of structure-activity relationships (Schneider et al., 2020; Wieder et al., 2020; Zhang et al., 2022; Fang et al., 2022; Wang et al., 2023; Chen et al., 2024). In this paradigm, atoms are usually treated as nodes and chemical bonds as edges, effectively encapsulating the connectivity that define molecular behaviors. However, it poses significant challenges due to intricate relationships among molecules and the limited chemical knowledge utilized during training.
@@
Figure 1: Multimodal Fusion with Relational Learning for Molecular Property Prediction (MMFRL). This figure shows our proposed idea about how to transfer the knowledge from other modalities and use fusion to improve the performance further. Unlike the general contrastive learning framework shown in Appendix Figure A.2, MMFRL does not need to define positive or negative pairs and is capable of learning continuous ordering from target similarity. In Early Fusion, a single Init GNN is created by combining all modality information during pretraining. In Intermediate and Late Fusion, each modality has its own initialized GNN.
@@
Contrastive Learning (CL) is often employed to study relationships among molecules. The primary focus within the domain of contrastive learning applied to molecular graphs centers on 2D-2D graph comparisons. Noteworthy representative examples: InfoGraph  (Sun et al., 2019) maximizes the mutual information between the representations of a graph and its substructures to guide the molecular representation learning; GraphCL  (You et al., 2020), MoCL  (Sun et al., 2021), and MolCLR  (Wang et al., 2022b) employs graph augmentation techniques to construct positive pairs; MoLR  (Wang et al., 2022a) establishes positive pairs with reactant-product relationships. In addition to 2D-2D graph contrastive learning, there are also noteworthy efforts exploring 2D-3D and 3D-3D contrastive learning in the field. 3DGCL  (Moon et al., 2023) is 3D-3D contrastive learning model, establishing positive pairs with conformers from the same molecules. GraphMVP  (Liu et al., 2022b), GeomGCL  (Li et al., 2022), and 3D Informax  (Stärk et al., 2022) proposes 2D–3D view contrastive learning approaches. To conclude, 2D-2D and 3D-3D comparisons are intra-modality contrastive learning, as only one graph encoder is employed in these studies. However, these approaches often focus on the motif and graph levels, leaving atom-level contrastive learning less explored. For example, consider Thalidomide: while the (R)- and (S)-enantiomers share the same topological graph and differ only at a single chiral center, their biological activities are drastically different—the (R)-enantiomer is effective in treating morning sickness, whereas the (S)-enantiomer causes severe birth defects. In other words, the (R)- and (S)-enantiomers are similar in terms of topological stucture but dissimilar in terms of biological activities. Thus, a more sophisticated approach is required to tackle these scenarios. A potential solution would be to use continuous metrics within a multi-view space, enabling a more comprehensive understanding of these complex molecular relationships.
There are multiple approaches for such similarity learning. One approach of them is instance-wise discrimination, which involves directly assessing the similarity between instances based on their latent representations or features.  (Wu et al., 2018b). Naive instance-wise discrimination relies on pairwise similarity, leading to the development of contrastive loss  (Hadsell et al., 2006). Although there are improved loss functions such as triplet loss  (Hoffer & Ailon, 2015), quadruplet loss  (Law et al., 2013), lifted structure loss  (Oh Song et al., 2016), N-pairs loss  (Sohn, 2016), and angular loss  (Wang et al., 2017), these methods still fall short in thoroughly capturing relationships among multiple instances simultaneously  (Wang et al., 2019). To address this limitation, a joint multi-similarity loss has been proposed, incorporating pair weighting for each pair to enhance instance-wise discrimination  (Wang et al., 2019; Zhang et al., 2021). However, these pair weightings requires the manual categorization of negative and positive pairs, as distinct weights are assigned to losses based on their categories. In this case, we can borrow the idea of Relational Learning  (Zheng et al., 2021) from computer vision by using different augmented views of the same instance tasks to similar features, while allowing for some variability. This approach captures the essential characteristics of the instance in a continuous scale, promoting relative consistency across the views without requiring them to be identical. By doing so, it enhances the model’s ability to generalize and recognize underlying patterns in the data.
@@
Besides, in order to enable a multi-view analysis from diverse sources is essential for improving molecule analysis, we can apply the Multi-Modality Fusion (Lahat et al., 2015; Khaleghi et al., 2013; Poria et al., 2015; Ramachandram & Taylor, 2017; Pawłowski et al., 2023; Manzoor et al., 2023; Priessner et al., 2024). It combines diverse heterogeneous data (e.g. text, images, graph) to create a more comprehensive understanding of complex scenarios . This approach leverages the strengths of each modality, potentially improving performance in tasks like sentiment analysis or medical diagnosis. Although challenging to implement due to the need to align different data streams, successful fusion can provide insights that that surpass those obtainable from individual modalities, advancing AI and data-driven decision-making. In particular, the way to fuse different modality should also depends on the dominance of each modality  (Pawłowski et al., 2023). However, when it comes to multimodal learning for molecules, we often encounter data availability and incompleteness issues. This raises a critical question: how can multimodal information be effectively leveraged for molecular property reasoning when such data is absent in downstream tasks? Recent studies have demonstrated the effectiveness of pretraining molecular Graph Neural Networks (GNNs) by integrating additional knowledge sources (Wang et al., 2021; 2022b; Liu et al., 2022a; Xu et al., 2023a). Building on this foundation, a promising solution is to pretrain multiple replicas of molecular GNNs, with each replica dedicated to learning from a specific modality. This approach allows downstream tasks to benefit from multimodal data that is not accessible during fine-tuning, ultimately improving representation learning.
@@
2 Results
2.1 The effectiveness of pre-training
We first illustrate the impact of pre-training initialization on performance on DMPNN  (Yang et al., 2019). As shown in Table 1, the average performance of pre-trained models outperform the non-pre-trained model in all tasks except for Clintox. The results of various downstream tasks indicate that different tasks may prefer different modalities. Notably, the model pre-trained with the NMR modality achieves the highest performance across three classification tasks. Similarly, the model pre-trained with the Image modality excels in three tasks, two of which are regression tasks related to solubility, aligning with findings from prior literature (Xu et al., 2023a). Additionally, the model pre-trained with The fingerprint method achieves the best performance in two tasks, including MUV, which has the largest dataset.
Max 93.4$\pm$2.7 93.2$\pm$2.7 68.1$\pm$1.5 87.7$\pm$6.5 83.3$\pm$1.1 82.9$\pm$3.1 86.0$\pm$1.0 71.1$\pm$0.8 0.761$\pm$ 0.068 1.437$\pm$ 0.134 0.537$\pm$ 0.005
2.2 Overall performance of MMFRL
As shown in Table 2 and Table 3, MMFRL demonstrates superior performance compared to all baseline models and the average performance of DMPNN pretrained with extra modalities across all 11 tasks evaluated in MoleculeNet. Results in Tables 5 and Table 6 demonstrates our great performance compared to the baseline models on the Dud-E (Mysinger et al., 2012) and LIT-PCBA (Tran-Nguyen et al., 2020) datasets. This robust performance highlights the effectiveness of our approach in leveraging multimodal data. In particular, while individual models pre-trained on other modalities for ClinTox fail to outperform the No-pretraining model, the fusion of these pre-trained models leads to improved performance. Besides, apart from Tox21 and Sider, the fusion models significantly enhances overall performance. In particular, the intermediate fusion model stands out by achieving the highest scores in seven distinct tasks, showcasing its ability to effectively combine features at a mid-level abstraction. the late fusion model achieves the top performance in two tasks. These results underscore the advantages of utilizing various fusion strategies in multimodal learning, further validating the efficacy of the MMFRL framework.
Data Set BBBP bace Sider Clintox HIV MUV Tox21 ToxCast
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Unimodality}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Max}}$ 93.4$\pm$2.7 93.2$\pm$2.7 68.1$\pm$1.5 87.7$\pm$6.5 83.3$\pm$1.1 82.9$\pm$3.1 86.0$\pm$1.0 71.1$\pm$0.8
Data Set ESOL FreeSolv Lipo
AttentiveFP 0.877$\pm$0.029 2.073$\pm$0.183 0.721$\pm$0.001
DMPNN 1.050$\pm$0.008 2.082$\pm$0.082 0.683$\pm$0.016
[…]
GEM 0.798$\pm$0.029 1.877$\pm$0.094 0.660$\pm$0.008
Uni-Mol 0.788$\pm$0.029 1.620$\pm$0.035 0.603$\pm$0.010
MolCLR 1.113$\pm$0.023 2.301$\pm$0.247 0.789$\pm$0.009
$\text{Unimodality}_{avg}$ 0.924$\pm$0.083 1.707$\pm$0.126 0.587$\pm$0.021
${Unimodality}_{Max}$ 0.761$\pm$ 0.068 1.437$\pm$ 0.134 0.537$\pm$ 0.005
$\text{MMFRL}_{early}$ 1.037$\pm$0.170 2.093$\pm$0.090 0.607$\pm$0.034
$\text{MMFRL}_{late}$ 0.763$\pm$0.035 1.741$\pm$0.191 0.525$\pm$0.018
Table 4: Study on the performances of $\text{MMFRL}_{Intermediate}$ with different contrastive loss functions: Contrastive Loss (CL) and Tripelet Loss (TL). The best results are denoted in bold. The first 8 datasets are classification tasks evaluated using ROC-AUC, while the last three are regression tasks evaluated using RMSE. Our model outperforms other loss functions in most of the datasets.
$\text{MMFRL}_{intermediate}$ 95.4$\pm$0.7 95.1$\pm$1.0 64.3$\pm$1.2 93.4$\pm$1.1 81.2$\pm$1.3 83.5$\pm$1.6 85.1$\pm$0.1 71.9$\pm$1.1 0.730$\pm$0.019 1.465$\pm$0.096 0.552$\pm$0.014
$\text{MMFRL}_{intermediateCL}$ 93.2$\pm$1.5 89.7$\pm$1.3 61.1$\pm$2.6 90.6$\pm$1.7 80.9$\pm$1.7 78.2$\pm$1.3 85.7$\pm$0.4 70.8$\pm$0.8 0.792$\pm$0.034 2.094$\pm$0.377 0.609$\pm$0.022
$\text{MMFRL}_{intermediateTL}$ 93.2$\pm$2.2 91.3$\pm$1.2 61.8$\pm$1.7 91.8$\pm$2.5 80.0$\pm$1.5 78.8$\pm$0.2 85.6$\pm$0.5 70.7$\pm$0.4 0.780$\pm$0.037 2.072$\pm$0.199 0.577$\pm$0.005
Table 5: Comparison of Average AUC-ROC across DUD-E dataset
Method Average AUC-ROC
COSP (Gao et al., 2022) 90.10
Graph CNN (Torng & Altman, 2019) 88.60
Drug VQA (Zheng et al., 2020) 97.20
AttentionSiteDTI (Yazdani-Jahromi et al., 2022) 97.10
$\text{DrugCLIP}_{FT}$ (Gao et al., 2023) 96.59
MMRFL-Intermediate (Ours) 98.32
MMRFL-Late (Ours) 96.78
Table 6: Overall performances (ROC-AUC) on classification downstream tasks for Lit-PCBA. The performance that are not by us is from the paper Cai et al. (2022)
Task NB SVM RF XGBoost DNN GCN GAT FP-GNN MMRFL_Int MMRFL_lat
ADRBZ 55.2 53.4 49.8 50.0 83.3 83.7 76.8 88.6 76.3 76.0
ALDH1 69.3 76.0 74.1 75.0 75.6 73.0 73.9 76.6 78.8 78.7
ESR1_ago 66.1 55.2 44.8 50.0 69.0 58.7 71.3 72.8 76.8 31.5
ESR1_ant 54.3 63.0 53.3 52.8 58.2 67.1 65.6 64.2 54.3 58.8
FEN1 87.6 87.7 65.7 88.8 90.1 89.7 88.8 88.9 90.5 85.4
GBA 70.9 77.8 59.9 83.0 77.7 73.5 77.6 75.1 77.3 75.3
IDH1 88.7 80.7 49.8 50.0 67.8 81.3 86.1 78.7 71.0 41.6
KAT2A 65.9 61.2 53.7 50.0 59.5 62.1 66.2 63.2 71.6 68.7
MAPK1 68.6 66.5 57.9 59.3 70.8 66.8 69.7 77.1 73.0 69.6
MTORC1 59.8 59.1 53.2 63.9 63.4 66.9 61.5 58.3 62.0 59.0
OPRK1 53.8 53.2 49.8 50.0 71.0 64.4 63.6 54.5 72.8 68.6
PKM2 68.4 75.3 58.1 73.7 71.9 63.6 72.4 73.2 77.2 71.4
PPARG 67.5 79.3 66.9 49.0 81.1 78.8 78.4 82.9 69.7 69.5
TP53 64.8 60.2 60.2 64.3 70.6 74.9 70.6 76.3 80.7 57.3
VDR 80.4 69.0 64.4 78.2 79.4 77.3 78.0 77.4 73.9 73.5
2.3 Analysis of the fusion effect
2.3.1 General Comparison among various Ways of Fusions
In contrast, Late Fusion enables each modality to be explored independently, maximizing the potential of individual modalities without interference from others. This separation allows for a thorough examination of each modality’s contribution. When certain modalities dominate the performance metrics, Late Fusion can capitalize on their strengths by effectively leveraging the most informative signals.. This approach is especially useful in scenarios where the dominance of specific modalities can be leveraged to enhance overall model performance.
In addition, we conduct an ablation study to evaluate the performance of our proposed loss functions against two traditional contrastive learning losses—Contrastive Loss and Triplet Loss—in the context of intermediate fusion. The experimental results as shown in Table 4 demonstrate that our proposed methods outperform the baseline approaches across the majority of tasks in the MoleculeNet dataset, thereby highlighting the superiority of our approach.
2.3.2 Explainability of Learnt Representations
To demonstrate the interpretability of the representations learned by the proposed fusion strategies, we present the post-hoc analysis results on two tasks, ESOL and Lipo, as case studies. The results showcase that the learnt representations can capture task-specific patterns and offer valuable insights for molecular design.
ESOL with Intermediate Fusion. As presented in Table 3, the intermediate fusion method  5.3.2 exhibits superior performance on the ESOL regression task for predicting solubility. To further analyze this performance, we employed t-SNE to reduce the dimensionality of the molecule embeddings from 300 to 2, resulting in a heatmap visualized in Figure 2. The embeddings derived from individual modalities prior to fusion do not display a clear pattern, showing no smooth transition from low to high solubility. In contrast, the embeddings by intermediate fusion reveal a distinct and smooth transition in solubility values: molecules with similar solubility cluster together, forming a gradient that extends from the bottom left (indicating lower solubility) to the upper center (representing higher solubility). This trend underscores the effectiveness of the intermediate fusion approach in accurately capturing the quantitative structure-activity relationships for aqueous solubility.
Lipo with Late Fusion. As detailed in Table 3, the Late Fusion method (described in Section  5.3.3) demonstrates superior performance on the Lipo regression task for predicting solubility in fats, oils, lipids, and non-polar solvents. According to Equation  11, the final prediction is determined by the respective coefficients ($w_{i}$) and predictions ($p_{i}$) from each modality.
Figure 4 shows the distributions of the coefficient values, predictions, and their products for each modality. Notably, the SMILES and Image modalities display a wide range of values, highlighting their potential to significantly influence the final predictions. This observation aligns with the strong performance achieved when pretraining using either of these two modalities, as shown in Table 1. In contrast, the $\text{NMR}_{\text{Peak}}$ values display a narrower range, indicating its role as a modifier for finer adjustments in the predictions. Furthermore, we observe that the contributions from $\text{NMR}_{\text{Spectrum}}$ and Fingerprint modalities are minimal, with their corresponding values approaching zero. This outcome highlights the advantages of the Late Fusion approach in effectively identifying and leveraging dominant modalities, thereby optimizing the overall predictive performance.
Figure 2: T-SNE visualization depicting the ESOL molecule embeddings for intermediate fusion in Section  5.3.2 alongside molecules within the highlighted region. Each point in the heatmap corresponds to the embeddings of respective molecules in ESOL, with color indicating solubility levels. Red denotes higher solubility, while blue indicates lower solubility. The embeddings derived from individual modalities prior to fusion do not display a clear pattern, the embeddings by intermediate fusion forms a gradient that extends from the bottom left (indicating lower solubility) to the upper center (representing higher solubility).
Substructure analysis with BACE. We explore the binding potential of positive inhibitor molecules targeting BACE and their associated key functional substructures, referred to as minimum positive subgraphs (MPS). To identify MPS, we employ a Monte Carlo Tree Search (MCTS) approach integrated into our BACE classification model, as implemented in RationalRL  (Jin et al., 2020). MCTS, being an iterative process, allows us to evaluate each candidate substructure for its binding potential with our model. Following the determination of MPSs, we categorize the original positive BACE molecules based on their respective MPSs. By computing the binding potential difference between the original molecule and its MPS, we can identify structural features that contribute to changes in binding affinity as shown in Figure  5.
In the case of the MPS 5 group, the binding score is heavily influenced by steric effects. The top three high-performing designs (5a–5c) all feature a flexible and compact alkylated pyrazole structure (colored green), which likely facilitates better accommodation within the binding pocket. In contrast, the three lowest-performing designs (5n–5p) incorporate a more rigid and bulky (trifluoromethoxy)benzene moiety (colored red), which may introduce steric hindrance and reduce binding efficiency. Additionally, the pyrazole ring contains two nitrogen atoms, offering more potential for hydrogen bonding interactions with the target protein, whereas the (trifluoromethoxy)benzene group has only one oxygen atom, limiting its capacity for such interactions. This comparison highlights the importance of both molecular flexibility and functional group composition in optimizing binding affinity.
Figure 5: The left sub-figure is the boxplot of the binding difference for the respective groups of molecules by the top 8 most frequent Minimum Positive Subgraph. The right sub-Figure showsthe detail strucutre of the 5th MPS.
2.3.3 Sensitivity Analysis
Choosing the most effective fusion strategy can be empirical. However, our results presented in Table 2,  3,  4, 5, and 6 provide strong evidence that our lightweight fusion strategy (early, intermediate, and late fusion) outperforms existing approaches in the literature. To guide the selection among these strategies, our intuition is as follows: if a modality is highly relevant to the downstream task, earlier fusion is likely to be more effective; otherwise, later fusion may be preferable.
To test this hypothesis, we performed a retrospective analysis to assess the sensitivity of downstream tasks to different fusion strategies. Since early fusion embeddings often lack the flexibility to adapt to individual samples, we excluded them from this analysis. Instead, we used pretrained encoders to extract embeddings for each modality and performed a simple linear regression between the embeddings and task labels. We then computed the Pearson correlation between the predicted values and the ground truth as a measure of each modality’s relevance.
For each dataset, we recorded the highest correlation across all modalities as the ”Top 1” score. We then concatenated the embeddings from all modalities and repeated the regression analysis. The improvement in correlation is reported as the ”Pearson Gain.” A higher Pearson Gain suggests that earlier fusion of multiple modalities is more beneficial. As shown in Table 7, datasets where intermediate fusion performs best generally exhibit higher Pearson Gain compared to late fusion, supporting our intuition. However, for ESOL and FreeSolv, the correlation from a single modality is already high, making them less suitable for this analysis.
Table 7: Person correlation of different modalities and chosen fusion strategies across datasets.
Dataset Smiles Image NMR FP Peak Top 1 Concat Pearson Gain Strategy
BBBP 0.757 0.765 0.733 0.746 0.484 0.765 0.935 0.170 Intermediate
Bace 0.752 0.759 0.757 0.772 0.474 0.772 0.957 0.185 Intermediate
Sider 0.598 0.579 0.578 0.592 0.403 0.597 0.973 0.376 Intermediate
Hiv 0.305 0.295 0.239 0.287 0.193 0.305 0.420 0.115 Late
MUV 0.159 0.159 0.141 0.147 0.050 0.159 0.315 0.156 Intermediate
Clintox 0.577 0.604 0.546 0.577 0.405 0.639 0.920 0.281 Intermediate
tox21 0.550 0.558 0.578 0.565 0.183 0.578 0.691 0.113 Intermediate
toxcast 0.587 0.590 0.523 0.577 0.333 0.590 0.908 0.318 Intermediate
Lipo 0.782 0.795 0.623 0.780 0.542 0.795 0.920 0.125 Late
ESOL 0.958 0.960 0.893 0.947 0.705 0.960 0.999 0.039 Intermediate
FreeSolv 0.982 0.980 0.915 0.977 0.768 0.982 0.999 0.017 Intermediate
3 Conclusion
4 Dataset
4.1 Selected Modalities for Target Similarity Calculation
Fingerprint: Fingerprints are binary vectors that represent molecular structures, capturing the presence or absence of particular substructures, fragments, or chemical features within a molecule. In particular, we utilize Morgan fingerprints, which are based on the Extended-Connectivity Fingerprints (ECFP) method introduced by Rogers and Hahn  (Rogers & Hahn, 2010). Specifically, we generate fingerprints using AllChem.GetMorganFingerprintAsBitVect(mol, 2), which corresponds to ECFP4 (radius = 2). Because ECFP4 is one the most effective and Interpretable Molecular Representations  (Zhong & Guan, 2023).
NMR (Nuclear Magnetic Resonance): NMR spectroscopy provides detailed insights into the chemical environment of atoms within a molecule  (Bunzel & Ralph, 2006). By analyzing the interactions of atomic nuclei with an applied magnetic field, NMR can reveal information about the structure, dynamics, and interactions of molecules, including the connectivity of atoms, functional groups, and conformational changes. In our experiments, $\text{NMR}_{\text{spectrum}}$ provides the information about the overal information of molecule while $\text{NMR}_{\text{peak}}$ provides the information about the individual atoms in the molecule.
4.2 Pre-training
4.3 Downstream tasks
Our model was trained on 11 drug discovery-related benchmarks sourced from MoleculeNet  (Wu et al., 2018a). Eight of them were classification tasks, including BBBP, BACE, SIDER, CLINTOX, HIV, MUV, TOX21, and ToxCas. The other three are regression tasks, including ESOL, Freesolv, and Lipo. Each dataset was divided into the train, validation, and test subsets in an 80%:10%:10% ratio using the scaffold splitter  (Halgren, 1996; Landrum, 2006) from Chemprop  (Yang et al., 2019; Heid et al., 2023). The scaffold splitter categorizes molecular data based on substructures, ensuring diverse structures in each subset. Molecules are partitioned into bins, with those exceeding half of the test set size assigned to training, promoting scaffold diversity in validation and test sets. Remaining bins are randomly allocated until reaching the desired set sizes, creating multiple scaffold splits for comprehensive evaluation.
The DUD-E (Directory of Useful Decoys: Enhanced) dataset  (Mysinger et al., 2012) is a widely used benchmark for virtual screening, containing 102 protein targets, thousands of active compounds, and carefully selected decoys that resemble actives in physico-chemical properties but differ topologically. In contrast, LIT-PCBA (Low-Throughput Informatics-Targeted PubChem BioAssay)  (Tran-Nguyen et al., 2020) offers a more realistic and challenging benchmark, derived from real experimental assays across 15 targets, with no artificial decoys and inherent data noise and imbalance. Together, they represent two ends of the spectrum in virtual screening evaluation—DUD-E with idealized conditions, and LIT-PCBA with real-world complexity. For the fine-tuning setting, We follow the same split and test approach as  (Gao et al., 2023) for DUD-E and  (Cai et al., 2022) for LIT-PCBA.
5 Methods
We first explain the preliminaries, and then our proposed modified metric in relational learning to facilitate smooth alignment between graph and referred unimodality. Then, we introduce approaches for integrating multi modalities at different stages of the learning process.
5.1 Molecular representation with DMPNN